theyadoreshayyy:

If sin y° = 7 divided by q and tan y° = 7 divided by r, what is the value of cos y°? cos y°= r divided by q cos y° = 7r cos y° = 7q cos y°= q divided by r

theyadoreshayyy:

jhonyy9:

how can you write tan using sin and cos ?

jhonyy9:

use this \[\tan x = \frac{ \sin x }{ \cos x }\]

theyadoreshayyy:

so how would i set it up ?

TyDaGreatest:

I know I'm not supposed to do this but there is no way I can explain it but the answer is A

jhonyy9:

@tydagreatest wrote:
I know I'm not supposed to do this but there is no way I can explain it but the answer is A
please give answer explained

TyDaGreatest:

ok

jhonyy9:

\[\sin y = \frac{ 7 }{ q }\] \[\tan y = \frac{ 7 }{ r }\]

jhonyy9:

@jhonyy9 wrote:
use this \[\tan x = \frac{ \sin x }{ \cos x }\]
replace these given details of sin y and tan y in this above wrote equation and calcule the cos y = ?

jhonyy9:

\[\tan y = \frac{ \sin y }{ \cos y }\]

jhonyy9:

\[\tan y = \frac{ 7 }{ r }\] \[\sin y = \frac{ 7 }{ q }\]